what angle is made when a clock reads 8 o'clock?
Answer by jim_thompson5910(35256)

Let the distance around the clock (from 12 to 12) be 12 units. The distance from 8 to 12 (in a clockwise fashion) is then 4 units (since 12 - 8 = 4)

So if the circumference of the clock is 12 units, and the hands subtend an arc that is 4 units, then the hands cut off 4/12 = 1/3 of the circumference.

There are 360 degrees in a circle. So the hands make an angle of 360*(1/3) = 360/3 = 120 degrees